The differential equation (DE) of y'' + (5^3 + sinx)^5 y' + y = cosx^3 is?


1
Expert's answer
2021-09-20T16:21:43-0400

"y'' + (5^3 + sinx)^5 y' + y = cosx^3"

Replace "y'" by "D" and y'' by "D^2" :

"D^2+ (5^3 + sinx)^5 D+ 1 = cosx^3\\
\Rightarrow D^2+ (5^3 + sinx)^5 D+ 1- cosx^3 =0\\
\Rightarrow D=\frac{-(5^3 + sinx)^5\pm \sqrt{((5^3 + sinx)^5)^2-4(1)(1-cosx^3)}}{2}"


Here the given differential equation is of order 2 and non-homogeneous.